# Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands

### v1

@@ Abstract
Scaling behavior of superfluid weight vs. band gap.—The dimensionless function $D_{s}/\Delta$ can be used to quantify the phase fluctuations; therefore, it is helpful to study the scaling behavior of $D_{s}/\Delta$ vs. $E_{g}/\Delta$, where the latter controls the crossover. In Fig. 4(a), we plot it at two fixed $\Delta$ values for the two-band model. Using Eq. 4, $D_{s}/\Delta$ can be expressed in terms of independent variables $\lambda\kappa$ and $\kappa$, where we find the change $\delta D_{s}^{\alpha}/\Delta$ has single scaling dependence on $\lambda\kappa$ and is insensitive to $\kappa$ value, i.e., all curves of the same type are parallel in Fig. 4(a). It is a general property of the weak interaction limit independent of the flat-band filling. The two contributions $D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}$ also show distinct scaling behaviors: $D_{s}^{\text{geo}}$ exhibits a weak logarithmic enhancement in the regime $E_{g}\gg\Delta$, but saturates around $E_{g}\sim\Delta$ due to the short-range nature of the crossover (see supplemental materials). In contrast, $D_{s}^{\text{conv}}$ only begins to grow after the dispersive bands become electron- or hole-doped at $E_{g}\sim\Delta$.
@@
Figure 4: (a) $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plot and (b) BKT $T_{c}$ for the two-band model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands, with $J=1$, $E_{g}=2m$ and $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}$. (a) shows the total ($D_{s}$, solid), geometric ($D_{s}^{\text{geo}}$, dashed) and conventional ($D_{s}^{\text{conv}}$, dot-dashed) contributions to the superfluid weight. In (b), $U_{A,B}$ are fixed at $0.99\varepsilon_{\Lambda}$ (blue) and $8.51\varepsilon_{\Lambda}$ (red), giving the uniform pairing $\Delta_{0}=0.01\varepsilon_{\Lambda}$ (blue) and $0.1\varepsilon_{\Lambda}$ (red) at $E_{g}=T=0$, respectively.
@@
BKT enhancement.— When the band gap closes, the ratio $D_{s}/\Delta>(J/4\pi)\ln(\varepsilon_{\Lambda}/2\Delta)$ diverges weakly as $\Delta\rightarrow 0$, indicating that for extremely weak interactions $T_{c}$ can be driven close to the mean-field temperature $T_{MF}$ by the singular gap closing. Calculating $D_{s}$ at finite temperatures and applying the Nelson-Kosterlitz criteria, we plot the BKT $T_{c}$ as a function of $E_{g}$ for the two-band model at two weak couplings in Fig. 4(b). Different from the $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plot, here we include the slight renormalization of $\Delta$ due to the change of density of states by the incoming dispersive band. We choose the superconducting gap at $T=E_{g}=0$, $\Delta_{0}$ as the energy unit since it approximately scales with the interaction $U$. This choice is more unbiased than the fixed unit $\varepsilon_{\Lambda}$ because then $T_{c}/\Delta_{0}$ contains the information of how efficiently the superconducting gap is converted to $T_{c}$. The $T_{c}$ enhancement for the large $\Delta_{0}$, large $E_{g}$ case (red curves) is larger than the other (blue curves) because its flat band has much weaker quantum geometry at $E_{g}/\Delta_{0}=5$. We also note that in the second phase of the gap-closing process ($E_{g}/\Delta_{0}<1$), the enhancement comes from $D_{s}^{\text{conv}}$ of the dispersive band; the geometric effect, although more singular, is canceled by the inter-band effect [16, 46].
@@
Strong interaction limit.—We now discuss the strong interaction limit ($\Delta\gg W_{d}$) of the gap closing effect on the superfluid weight by analyzing case (ii) of the three-band model. We make a complete comparison between a singular and non-singular gap closing within the same model by choosing $c(m)=m$ for the former and $c(m)=c_{0}$ for the latter, where $E_{g}=c_{0}$ then is independent of $m$. For convenience, we fix $m$ to be $m_{0}=0.42\varepsilon_{\Lambda}$ for the non-singular model. At this value, the three orbitals have equal weight in all three bands, so the uniform pairing condition is satisfied. A non-singular band gap closing has the general property that the quantum metric $g_{\mu\nu}$ may vary with k and the band gap $E_{g}$, but never diverges when $E_{g}$ completely closes. It implies that the function $g_{\mu\nu}(\textbf{k},E_{g})$ is bounded, which justifies the choice of $c(m)=c_{0}$ for the non-singular case. For the flattened model, the conventional superfluid weight vanishes, therefore we only discuss the geometric superfluid weight below.
@@
Since the width of dispersive bands is zero, the nature of crossover for $D_{s}^{\text{geo}}$ is also changed, with a crossover function $\chi_{f}(\lambda\kappa)=1-(1+\lambda^{2}\kappa^{2})^{-1/2}$ instead of Eq. 5. $\chi_{f}$ depends on the single parameter $\lambda\kappa=E_{g}/\Delta$, so $D_{s}^{\text{geo}}\propto\Delta\chi_{f}$ exhibits a $\Delta$ to $1/\Delta$ crossover in the regime $E_{g}\sim\Delta$. This is expected for the strong interaction case when the interaction scale includes all the bands and the superfluid weight vanishes as $1/\Delta$. What distinguishes the singular gap closing is the $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plot, as shown by the solid lines in Fig. 5(a), where a maximum is achieved at $E_{g}/\Delta\sim 1$ and then decreases to zero at $E_{g}=0$, in contrast to a monotonic decrease for the non-singular gap closing (dashed curve). For the strong interaction case, $\Delta$ renormalizes with $E_{g}$ more strongly, leading to a competition between $\Delta$ and $D_{s}$ as $E_{g}$ changes, which determines $E_{g}\sim\Delta$ as the optimal condition for $T_{c}$ of 2D superconductors. In Fig. 5(b), we plot the BKT transition temperature $T_{c}$ for both the singular and non-singular cases. The singular case has a sharper peak at $E_{g}\sim\Delta$ than the non-singular case since in the regime $E_{g}>\Delta$, both $\Delta$ and $D_{s}$ enhance as $E_{g}$ decreases.
@@
Discussion.—Nontrivial band geometry is an essential feature of flat bands that cannot originate from spatially decoupled atomic orbitals. We show that the interplay of the band gap $E_{g}$ and superconducting gap $\Delta$ determines the crossover behavior of the superfluid weight in singular flat bands. Similar behavior is expected for systems with concentrated quantum metric that scales as $g_{\mu\nu}(\textbf{k})\sim k^{-2}$ [47]. In finite-hopping tight-binding models, singular flat bands originating from quantum interference of Bloch phases [30, 32] are also characterized by non-analytic Bloch eigenstates [31, 41]. In this case, for the two-band model, the flat band acquires a dispersion upon adding a mass $m$, with the flat bandwidth $W_{f}$ proportional to $m$, whereas it remains dispersionless in the three-band model. For these systems with $W_{f}\sim m$, the geometric superfluid weight will be approximately $\Delta\ln(W_{d}/\text{max}\{W_{f},\Delta\})$ during the gap closing. Our analysis also reveals that the supercurrent and BKT transition temperature is enhanced and optimized by gap closing in a singular flat band system, paving the road for searching for high-$T_{c}$ flat-band superconductors.
@@
Acknowledgement.—This work was supported by the award number DE-SC0022178 from the Department of Energy (Y.B.), the UNR/VPRI Startup funds PG19012 (Y.B & G.J.) and the Research Council of Finland (former Academy of Finland) under project number 339313 and by Jane and Aatos Erkko Foundation, Keele Foundation, and Magnus Ehrnrooth Foundation as part of the SuperC collaboration (P.T & G.J.).
@@
Supplemental Materials for “Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands”
@@ I S1: Flat band touching in short-range hopping models
Finally, we would like to remark the two-band model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands in the main text. In Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands, the $\hat{\sigma}_{0}$ term coefficient cannot be expressed as a polynomial of $k_{x},k_{y}$. Therefore, it introduces infinite-range hoppings, which makes the valence band perfectly flat. However, when the chirality index $J$ is even, the $m=0$ case of Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands can be realized in finite-range models, as the coefficient of $\hat{\sigma}_{0}$ becomes $\zeta_{J}k^{J}$ which is a polynomial of $k_{x},k_{y}$ ($J=2$ corresponds to Kagome and Mielke lattice singular flat bands [30, 35, 31, 32]). Similarly, one can check that for the three-band model Eq. 3 the unflattened case (i) can always be realized in finite-range models, whereas the flattened case (ii) cannot.
@@ II S2: Superfluid Weight Calculation for Singular Two-Band Models
$+\sum_{l,\textbf{k}}\tanh\frac{\beta E_{l\textbf{k}}}{2}\bigg{\{}\frac{4\Delta^{2}}{E_{l\textbf{k}}}g^{l}_{\mu\nu}(\textbf{k})+\sum_{l^{\prime}\neq l}8\Delta^{2}\bigg{[}\frac{p^{(+)}_{ll^{\prime}}(\textbf{k})}{E_{l\textbf{k}}+E_{l^{\prime}\textbf{k}}}+\frac{p^{(-)}_{ll^{\prime}}(\textbf{k})}{E_{l\textbf{k}}-E_{l^{\prime}\textbf{k}}}\bigg{]}g^{ll^{\prime}}_{\mu\nu}(\textbf{k})\bigg{\}},$ (8)
@@
$D_{s,\mu\nu}(T=0)=\sum_{l,\textbf{k}}\bigg{[}\frac{\Delta^{2}}{E_{l\textbf{k}}^{3}}\partial_{\mu}\xi_{l\textbf{k}}\partial_{\nu}\xi_{l\textbf{k}}+\frac{4\Delta^{2}}{E_{l\textbf{k}}}g^{l}_{\mu\nu}(\textbf{k})\bigg{]}+\sum_{l^{\prime}>l,\textbf{k}}\frac{16\Delta^{2}p^{(+)}_{ll^{\prime}}(\textbf{k})}{E_{l\textbf{k}}+E_{l^{\prime}\textbf{k}}}g^{ll^{\prime}}_{\mu\nu}(\textbf{k}).$ (9)
They are consistent with the general expression in Ref. [5, 18]. Here $l,l^{\prime}$ are band indices, $\Delta$ is the uniform pairing order parameter either at finite temperature $T=1/\beta$ or $T=0$, and $E_{l\textbf{k}}=\sqrt{\xi_{l\textbf{k}}^{2}+\Delta^{2}}$ is the quasiparticle energy, with $\xi_{l\textbf{k}}=\varepsilon_{l\textbf{k}}-\mu$. The first and second line of Eq. II correspond to the conventional and geometric contributions to superfluid weight, respectively.
$p^{(\pm)}_{ll^{\prime}}(\textbf{k})=\frac{1}{2}\bigg{(}1\pm\frac{\xi_{l\textbf{k}}\xi_{l^{\prime}\textbf{k}}+\Delta^{2}}{E_{l\textbf{k}}E_{l^{\prime}\textbf{k}}}\bigg{)},\,\,\,l\neq l^{\prime},$ (10)
@@
$g^{ll^{\prime}}_{\mu\nu}(\textbf{k})=-\text{Re}\langle\partial_{\mu}u_{l\textbf{k}}|u_{l^{\prime}\textbf{k}}\rangle\langle u_{l^{\prime}\textbf{k}}|\partial_{\nu}u_{l\textbf{k}}\rangle,\,\,\,l\neq l^{\prime}.$ (11)
@@
For the two-band model Eq. Superfluid Weight Crossover and Critical Temperature Enhancement in Singular Flat Bands, since the valence band remains flat when the band gap is open, the conventional contribution of superfluid weight only comes from the conduction band. At $T=0$, it can be expressed as
$D_{s}^{\text{conv}}=\sum_{\textbf{k}}\frac{\Delta^{2}}{E_{c{\bf k}}^{3}}(\partial_{x}\varepsilon_{c{\bf k}})^{2},$ (12)
with $\varepsilon_{c{\bf k}}=2\sqrt{\zeta_{J}^{2}k^{2J}+m^{2}}$ the conduction band energy. Evaluating Eq. 12 with a momentum cutoff $k_{\Lambda}$ gives
$D_{s}^{\text{conv}}=\frac{J\Delta}{4\pi}\bigg{[}\lambda^{2}\kappa^{2}\bigg{(}\ln\frac{1+a_{\kappa}}{1+a_{\lambda\kappa}}+\ln\lambda\bigg{)}+a_{\lambda\kappa}-\frac{a_{\lambda\kappa}}{a_{\kappa}}\bigg{]},$ (13)
@@
$D_{s}^{\text{geo}}=4\Delta\sum_{{\bf k}}\bigg{(}1-\frac{\Delta}{E_{c{\bf k}}}\bigg{)}g_{xx}({\bf k}).$ (14)
Eq. 14 can be further evaluated using the quantum metric expression Eq. 2. We find the valence and conduction band geometric superfluid weights are
$D^{\text{geo},v}_{s}=\frac{J\Delta}{8\pi}(1-\lambda^{2}-2\ln\lambda),$ (15)
$D_{s}^{\text{geo},c}=\frac{J\Delta}{8\pi}\bigg{[}(\lambda^{2}\kappa^{2}-2)\bigg{(}\ln\frac{1+a_{\lambda\kappa}}{1+a_{\kappa}}+\ln\lambda\bigg{)}+a_{\lambda\kappa}-\lambda^{2}a_{\kappa}\bigg{]}.$ (16)
@@
$h(\textbf{k})=v(k_{x}\hat{\sigma}_{x}+k_{y}\hat{\sigma}_{y}),$ (17)
@@
$g_{\mu\nu}(\textbf{k})=\frac{k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu}}{4k^{4}}.$ (18)
@@
Aligning the chemical potential with the Dirac point, $\mu=0$, the problem has two energy scales $vk_{\Lambda}$ and $\Delta$, with $k_{\Lambda}$ the cutoff momentum, from which only one dimensionless parameter $\kappa=vk_{\Lambda}/\Delta$ can be defined. The zero-temperature superfluid weight can be found using Eq. 9, giving the conventional and geometric superfluid weight
$D_{s}^{\text{conv}}=\frac{\Delta}{\pi}\bigg{(}1-\frac{1}{a_{\kappa}}\bigg{)}=2D_{s}^{\text{geo}},$ (19)

### v2

@@ Abstract
Scaling behavior of superfluid weight vs band gap.—The dimensionless function $D_{s}/\Delta$ can be used to quantify the phase fluctuations; therefore, it is helpful to study the scaling behavior of $D_{s}/\Delta$ vs $E_{g}/\Delta$, where the latter controls the cross-over. In Fig. 4(a), we plot it at two fixed $\Delta$ values for the two-band model for $J=1$. Using Eq. 5, $D_{s}/\Delta$ can be expressed in terms of independent variables $\lambda\kappa$ and $\kappa$, where we find the change $\delta D_{s}^{\alpha}/\Delta$ has single scaling dependence on $\lambda\kappa$ and is insensitive to $\kappa$ value, i.e., all curves of the same type are parallel in Fig. 4(a). This is a general property of the weak interaction limit and is independent of the flat-band filling. The two contributions $D_{s}^{\text{geo}}$ and $D_{s}^{\text{conv}}$ also show distinct scaling behaviors: $D_{s}^{\text{geo}}$ exhibits a weak logarithmic enhancement in the regime $E_{g}\gg\Delta$, but saturates around $E_{g}\sim\Delta$ due to the short-range nature of the cross-over (see supplemental materials Sec. S4). In contrast, $D_{s}^{\text{conv}}$ only begins to grow after the dispersive bands become electron- or hole-doped at $E_{g}\sim\Delta$.
@@
Figure 4: (a) $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plot and (b) BKT $T_{c}$ for the two-band model Eq. Superfluid weight cross-over and critical temperature enhancement in singular flat bands, with $J=1$, $E_{g}=2m$ and $\varepsilon_{\Lambda}=2\zeta_{J}k_{\Lambda}^{J}$. (a) shows the total ($D_{s}$, solid), geometric ($D_{s}^{\text{geo}}$, dashed) and conventional ($D_{s}^{\text{conv}}$, dot-dashed) contributions to the superfluid weight. In (b), $U_{A,B}$ are fixed at $0.99\varepsilon_{\Lambda}$ (blue) and $8.51\varepsilon_{\Lambda}$ (red), giving the uniform pairing $\Delta_{0}=0.01\varepsilon_{\Lambda}$ (blue) and $0.1\varepsilon_{\Lambda}$ (red) at $E_{g}=T=0$, respectively.
@@
BKT enhancement.—When the band gap closes, the ratio $D_{s}/\Delta>(J/4\pi)\ln(\varepsilon_{\Lambda}/2\Delta)$ diverges weakly as $\Delta\rightarrow 0$, indicating that for extremely weak interactions $T_{c}$ can be driven close to the mean-field temperature $T_{MF}$ by the singular gap closing. Calculating $D_{s}$ at finite temperatures and applying the Nelson-Kosterlitz criteria, we plot the BKT $T_{c}$ as a function of $E_{g}$ for the two-band model for $J=1$ at two weak couplings in Fig. 4(b). Different from the $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plot, here we include the slight renormalization of $\Delta$ due to the change of density of states by the incoming dispersive band. We choose the superconducting gap at $T=E_{g}=0$, $\Delta_{0}$ as the energy unit since it approximately scales with the interaction $U$. This choice is more unbiased than the fixed unit $\varepsilon_{\Lambda}$ because then $T_{c}/\Delta_{0}$ contains the information of how efficiently the superconducting gap is converted to $T_{c}$. The $T_{c}$ enhancement for the large $\Delta_{0}$, large $E_{g}$ case (red curve) is larger than the other (blue curve) because its flat band has much weaker quantum geometry at $E_{g}/\Delta_{0}=5$. We also note that in the second phase of the gap closing process ($E_{g}/\Delta_{0}<1$), the enhancement comes from $D_{s}^{\text{conv}}$ of the dispersive band; the geometric effect, although more singular, is canceled by the inter-band effect [20, 34].
@@
Strong interaction limit.—We now discuss the strong interaction limit ($\Delta\gg W_{d}$) of the gap closing effect on the superfluid weight by analyzing case (ii) of the three-band model. We compare gap closing of a singular and non-singular flat band within the same model at $J=1$, by choosing $c(m)=m$ for the former and $c(m)=c_{0}$ for the latter, where $E_{g}=c_{0}$ then is independent of $m$. For convenience, we also fix $m$ to be $m_{0}=0.42\varepsilon_{\Lambda}$ for the non-singular model. At this value, the three orbitals have equal weight in all three bands, so the uniform pairing condition is satisfied. A non-singular band gap closing has the general property that the quantum metric $g_{\mu\nu}$ may vary with k and the band gap $E_{g}$, but never diverges when $E_{g}$ completely closes. It implies that the function $g_{\mu\nu}(\textbf{k},E_{g})$ is bounded, which justifies the choice of $c(m)=c_{0}$ for the non-singular case. For the flattened model, the conventional superfluid weight vanishes, therefore we only discuss the geometric superfluid weight below.
@@
Since the width of the dispersive bands is zero, the nature of cross-over for $D_{s}^{\text{geo}}$ is also changed, with a cross-over function $\chi_{f}(\lambda\kappa)=1-(1+\lambda^{2}\kappa^{2})^{-1/2}$ instead of Eq. 6 ($\chi_{f}$ can be derived from Eq. 4 by taking the strong interaction limit). $\chi_{f}$ depends on the single parameter $\lambda\kappa=E_{g}/\Delta$, so $D_{s}^{\text{geo}}\propto\Delta\chi_{f}$ exhibits a $\propto\Delta$ to $1/\Delta$ cross-over in the regime $E_{g}\sim\Delta$. This is expected for the strong interaction case when the interaction scale includes all the bands and the superfluid weight vanishes as $1/\Delta$. What distinguishes the singular gap closing is the $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plot, as shown by the solid curves in Fig. 5(a), where a maximum is achieved at $E_{g}/\Delta\sim 1$ and then decreases to zero at $E_{g}=0$, in contrast to a monotonic decrease for the non-singular gap closing (dashed curve). In the strong interaction case, $\Delta$ is enhanced by the gap closing more strongly than in the weak interaction case, leading to a competition between $\Delta$ and $D_{s}$ as $E_{g}$ changes, which determines $E_{g}\sim\Delta$ as the optimal condition for $T_{c}$ of 2D superconductors. In Fig. 5(b), we plot the BKT transition temperature $T_{c}$ for both the singular and non-singular cases. The singular case has a sharp peak at $E_{g}\sim\Delta$, since in the regime $E_{g}>\Delta$, both $\Delta$ and $D_{s}$ increase as $E_{g}$ decreases. On the contrary, although $\Delta$ is enhanced, $T_{c}$ of the non-singular case is almost unchanged at $E_{g}>\Delta$ due to the decrease of $D_{s}/\Delta$. As $E_{g}/\Delta$ gets smaller than 1, $T_{c}$ of both cases decreases rapidly due to the inter-band cancellation effect for $D_{s}$ [20, 34]. This cancellation ($D_{s}\rightarrow 0$, as $E_{g}\rightarrow 0$) behavior is due to $E_{c,\textbf{k}}\rightarrow\Delta$ in the second term of Eq. 4 when the dispersive bandwidth becomes smaller than $\Delta$ over the entire Brillouin zone.
Results for large chirality $J$.—As $J$ increases, the band geometry gets stronger; the dispersive bands exhibit a $k^{J}$-dispersion when the gap closes  [^2] These $J$-chiral models resemble the effective model of $J$-layer rhombohedral stacked multi-layer with only nearest-neighbor inter-layer hoppings. In general, remote hopping effects lead to the trigonal warping of bands and instability for large values of $J$, resulting in smaller values of $J$-chiral models at neighboring points in the BZ.. As a result, the $D_{s}/\Delta$ axis of the $D_{s}/\Delta$ vs $E_{g}/\Delta$ scaling plots for all models (Fig. 4(a) and 5(a)) will scale with $J$, like the two-band case shown in Table 1, leading to the possibility of $D_{s}/\Delta\gg 1$ for large $J$. Then, the phase fluctuation is weak ($T_{c}/T_{MF}\approx 1$), and $T_{c}$ is limited by the superconducting gap $\Delta$ rather than $D_{s}$. In other words, too strong band geometry and dispersion weaken the role of $D_{s}$ enhancement through singular band gap closing. For completeness, we present the results for the $J=6$ case in the supplementary materials Sec. S8. For large values of $J$ with strong interactions, non-singular gap closing has the same sharp $T_{c}$ enhancement peak as the singular gap closing since both models experience the same enhancement of $\Delta$ due to the change in the density of states (see Fig. S8(b) in the supplementary materials, cf. Fig. 5(b)).
@@
Discussion.—The mechanism to enhance the critical temperature in quasi-flat band superconductors is not unique. Recent studies have found that the conventional superfluid weight in quasi-flat bands can be enhanced by interband pair-exchange interactions [52]. However, nontrivial band geometry is a fundamental property of flat bands that cannot originate from spatially decoupled atomic orbitals. We showed that the interplay between the band gap $E_{g}$ and superconducting gap $\Delta$ determines the cross-over behavior of the superfluid weight in singular flat bands, which is purely geometric in origin. Similar behavior is expected for systems with concentrated quantum metric that scales as $g_{\mu\nu}(\textbf{k})\propto k^{-2}$ [53]. In finite-hopping tight-binding models, singular flat bands originating from quantum interference of Bloch phases [35, 37] are also characterized by non-analytic Bloch eigenstates [36, 46]. In this case, for the two-band model, the flat band acquires a dispersion upon adding a mass $m$, with the flat bandwidth $W_{f}$ proportional to $m$, whereas it remains dispersionless in the three-band model. For these systems with $W_{f}\sim m$, the geometric superfluid weight will be approximately $\Delta\ln(W_{d}/\text{max}\{W_{f},\Delta\})$ during the gap closing. Our analysis also reveals that the supercurrent and BKT transition temperature is enhanced and optimized by gap closing in a singular flat band system, paving the road for searching for high-$T_{c}$ flat-band superconductors.
@@
Acknowledgement.—The authors would like to thank Andrew J. Millis and Grigory E. Volovik for helpful discussions. This work was supported by the award number DE-SC0022178 from the Department of Energy (Y.B.), the UNR/VPRI Startup funds PG19012 (Y.B & G.J.) and the Research Council of Finland (former Academy of Finland) under project number 339313 and by Jane and Aatos Erkko Foundation, Keele Foundation, and Magnus Ehrnrooth Foundation as part of the SuperC collaboration (P.T. & G.J.).
@@
Supplementary Information for “Superfluid weight cross-over and critical temperature enhancement in singular flat bands”
@@ I S1: Flat band touching in short-range hopping models
Finally, we would like to remark on the two-band model Eq. 1 in the main text (MT). In Eq. 1, the $\hat{\sigma}_{0}$ term coefficient cannot be expressed as a polynomial of $k_{x},k_{y}$. Therefore, it introduces infinite-range hoppings, which makes the valence band perfectly flat. However, when the chirality index $J$ is even, the $m=0$ case of Eq. 1 can be realized in finite-range models, as the coefficient of $\hat{\sigma}_{0}$ becomes $\zeta_{J}k^{J}$ which is a polynomial of $k_{x},k_{y}$ ($J=2$ corresponds to Kagome and Mielke lattice singular flat bands [35, 40, 36, 37]). Similarly, one can check that for the three-band model Eq. 3 the unflattened case (i) can always be realized in finite-range models, whereas the flattened case (ii) cannot be realized with a local Hamiltonian.
@@ II S2: Superfluid Weight Calculation for Singular Two-Band Models
$+\sum_{l,\textbf{k}}\tanh\frac{\beta E_{l\textbf{k}}}{2}\bigg{\{}\frac{4\Delta^{2}}{E_{l\textbf{k}}}g^{l}_{\mu\nu}(\textbf{k})+\sum_{l^{\prime}\neq l}8\Delta^{2}\bigg{[}\frac{p^{(+)}_{ll^{\prime}}(\textbf{k})}{E_{l\textbf{k}}+E_{l^{\prime}\textbf{k}}}+\frac{p^{(-)}_{ll^{\prime}}(\textbf{k})}{E_{l\textbf{k}}-E_{l^{\prime}\textbf{k}}}\bigg{]}g^{ll^{\prime}}_{\mu\nu}(\textbf{k})\bigg{\}},$ (S1)
@@
$D_{s,\mu\nu}(T=0)=\sum_{l,\textbf{k}}\bigg{[}\frac{\Delta^{2}}{E_{l\textbf{k}}^{3}}\partial_{\mu}\xi_{l\textbf{k}}\partial_{\nu}\xi_{l\textbf{k}}+\frac{4\Delta^{2}}{E_{l\textbf{k}}}g^{l}_{\mu\nu}(\textbf{k})\bigg{]}+\sum_{l^{\prime}>l,\textbf{k}}\frac{16\Delta^{2}p^{(+)}_{ll^{\prime}}(\textbf{k})}{E_{l\textbf{k}}+E_{l^{\prime}\textbf{k}}}g^{ll^{\prime}}_{\mu\nu}(\textbf{k}).$ (S2)
This is consistent with the general expression in Ref. [6, 22]. Here $l,l^{\prime}$ are band indices, $\Delta$ is the uniform pairing order parameter either at finite temperature $T=1/\beta$ or $T=0$, and $E_{l\textbf{k}}=\sqrt{\xi_{l\textbf{k}}^{2}+\Delta^{2}}$ is the quasiparticle energy, with $\xi_{l\textbf{k}}=\varepsilon_{l\textbf{k}}-\mu$. The first and second line of Eq. II correspond to the conventional and geometric contributions to superfluid weight, respectively.
$p^{(\pm)}_{ll^{\prime}}(\textbf{k})=\frac{1}{2}\bigg{(}1\pm\frac{\xi_{l\textbf{k}}\xi_{l^{\prime}\textbf{k}}+\Delta^{2}}{E_{l\textbf{k}}E_{l^{\prime}\textbf{k}}}\bigg{)},\,\,\,l\neq l^{\prime},$ (S3)
@@
$g^{ll^{\prime}}_{\mu\nu}(\textbf{k})=-\text{Re}\langle\partial_{\mu}u_{l\textbf{k}}|u_{l^{\prime}\textbf{k}}\rangle\langle u_{l^{\prime}\textbf{k}}|\partial_{\nu}u_{l\textbf{k}}\rangle,\,\,\,l\neq l^{\prime}.$ (S4)
@@
For the two-band model Eq. 1 in MT, since the valence band remains flat when the band gap is open, the conventional contribution of superfluid weight only comes from the conduction band. At $T=0$, it can be expressed as
$D_{s}^{\text{conv}}=\sum_{\textbf{k}}\frac{\Delta^{2}}{E_{c{\bf k}}^{3}}(\partial_{x}\varepsilon_{c{\bf k}})^{2},$ (S5)
with $\varepsilon_{c{\bf k}}=2\sqrt{\zeta_{J}^{2}k^{2J}+m^{2}}$ the conduction band energy. Evaluating Eq. (S5) with a momentum cutoff $k_{\Lambda}$ gives
$D_{s}^{\text{conv}}=\frac{J\Delta}{4\pi}\bigg{[}\lambda^{2}\kappa^{2}\bigg{(}\ln\frac{1+a_{\kappa}}{1+a_{\lambda\kappa}}+\ln\lambda\bigg{)}+a_{\lambda\kappa}-\frac{a_{\lambda\kappa}}{a_{\kappa}}\bigg{]},$ (S6)
@@
$D_{s}^{\text{geo}}=4\Delta\sum_{{\bf k}}\bigg{(}1-\frac{\Delta}{E_{c{\bf k}}}\bigg{)}g_{xx}({\bf k}).$ (S7)
Eq. (S7) can be evaluated using the quantum metric expression Eq. 2 in MT. We find the valence and conduction band geometric superfluid weights are
$D^{\text{geo},v}_{s}=\frac{J\Delta}{8\pi}(1-\lambda^{2}-2\ln\lambda),$ (S8)
$D_{s}^{\text{geo},c}=\frac{J\Delta}{8\pi}\bigg{[}(\lambda^{2}\kappa^{2}-2)\bigg{(}\ln\frac{1+a_{\lambda\kappa}}{1+a_{\kappa}}+\ln\lambda\bigg{)}+a_{\lambda\kappa}-\lambda^{2}a_{\kappa}\bigg{]}.$ (S9)
@@
$h(\textbf{k})=v(k_{x}\hat{\sigma}_{x}+k_{y}\hat{\sigma}_{y}),$ (S10)
@@
$g_{\mu\nu}(\textbf{k})=\frac{k^{2}\delta_{\mu\nu}-k_{\mu}k_{\nu}}{4k^{4}}.$ (S11)
@@
Aligning the chemical potential with the Dirac point, $\mu=0$, the problem has two energy scales $vk_{\Lambda}$ and $\Delta$, with $k_{\Lambda}$ the cutoff momentum, from which only one dimensionless parameter $\kappa=vk_{\Lambda}/\Delta$ can be defined. The zero-temperature superfluid weight can be found using Eq. (S2), giving the conventional and geometric superfluid weight
$D_{s}^{\text{conv}}=\frac{\Delta}{\pi}\bigg{(}1-\frac{1}{a_{\kappa}}\bigg{)}=2D_{s}^{\text{geo}},$ (S12)